OpenStudy (anonymous):

help 40 ÷ [24 – 4 • (2 + 3)]

OpenStudy (hyper):

Still need help?

OpenStudy (texaschic101):

40 / [ 24 - 4 * (2 + 3) ] = 40 / [ 24 - 4 * 5 ] = 40 / [ 24 - 20 ] = 40 / 4 = 10

OpenStudy (anonymous):

thanks

OpenStudy (texaschic101):

no problem :)